### v1

# Beyond the Request: Harnessing HTTP Response Headers for Cross-Browser Web Tracker Classification in an Imbalanced Setting

Abstract.
The World Wide Web’s connectivity is greatly attributed to the HTTP protocol, with HTTP messages offering informative header fields that appeal to disciplines like web security and privacy, especially concerning web tracking. Despite existing research employing HTTP/S request messages to identify web trackers, HTTP/S response headers are often overlooked. This study endeavors to design effective machine learning classifiers for web tracker detection using HTTP/S response headers. Data from the Chrome, Firefox, and Brave browsers, obtained through the traffic monitoring browser extension T.EX, serves as our data set. Eleven supervised models were trained on Chrome data and tested across all browsers. The results demonstrated high accuracy, F1-score, precision, recall, and minimal log-loss error for Chrome and Firefox, but subpar performance on Brave, potentially due to its distinct data distribution and feature set. The research suggests that these classifiers are viable for detecting web trackers in Chrome and Firefox. However, real-world application testing remains pending, and the distinction between tracker types and broader label sources could be explored in future studies.
Web Privacy, Web Tracking, Web Privacy Measurement, HTTP/S Response Header, Machine Learning
journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
1. Introduction
The rapid evolution of the Web has cemented its role as an integral component of modern life. Through desktop and mobile browsers, users access a vast array of information and services. However, every interaction, from content consumption to mere site visitation, reveals information about the habits and preferences of users, which in turn has implications for personal privacy and online security.
These privacy concerns are highlighted by the ubiquity of behavioural targeting. By analyzing a user’s browsing history, a detailed profile can be constructed, facilitating the delivery of tailored advertisements and content. This profiling is made possible by web tracking that employs sophisticated techniques to track users across the Web. Given the increasing concerns about such practices (Melicher et al., 2016; Chanchary and Chiasson, 2015), there is a growing emphasis on research in web privacy.
An established countermeasure against unsolicited HTTP communications is the use of filter lists (Traverso et al., 2017), such as EasyList (eas, 2023a) and EasyPrivacy (eas, 2023b). However, their dependency on manual curation makes them vulnerable to oversights and rapid obsolescence (Hashmi et al., 2019; Alrizah et al., 2019; Fouad et al., 2018). Moreover, some trackers employ strategies like domain masking to evade these lists (Chen et al., 2021).
In an attempt to mitigate these shortcomings, recent efforts have explored the potential of Machine Learning (ML) in this domain. Various classifiers, some achieving accuracy metrics in the 80th percentile (Cozza et al., 2020; Iqbal et al., 2021), have been developed using diverse data sources such as HTTP headers and JavaScript calls. Yet, a majority of these focused primarily on HTTP requests, potentially neglecting other informative data streams. Additionally, the predominant reliance on Firefox for data collection may limit the comprehensiveness of these studies.
Notably, there seems to be a gap in literature regarding the use of HTTP response headers as the sole data source for tracker classification across diverse browsers, despite studies in the related field of online security highlighting the efficacy of this approach (McGahagan et al., 2019). This study intends to bridge this gap. We strive to develop classifiers adept at web tracker detection, while confronting challenges such as data sparsity, high dimensionality, and class imbalance.
Our key contributions are:
  1. (1)
    We conduct a rigorous study of HTTP response headers across multiple browsers, focusing on their potential for web tracker detection within the Tranco top 10k websites. Furthermore, we evaluate dataset similarities to discern differences in HTTP response headers across the assessed browsers.
  2. (2)
    We present a semi-automated ML pipeline which facilitates the training of eleven classifiers tailed for web tracker detection. Our findings indicate that selected tree and gradient boosting models, when applied to a subset of binary-encoded HTTP/S headers, are capable of achieving ROC-AUC, AUPCR, and F1 scores exceeding $0.9$.
  3. (3)
    We conduct a comprehensive and multi-faceted evaluation of these classifiers, which includes longitudinal studies to gauge potential concept-drift and evaluates the cross-browser performance of classifiers trained solely on traffic data collected within the Chrome browser. Notably, our evaluation incorporates thirteen metrics, offering a holistic view of the binary classification challenge.
  4. (4)
    We assess calibration effects on final classifier performance — a consideration seldom addressed in prior research. This allows us to discern if models trained solely on data from one browser can be effectively deployed on data from alternate browsers without the necessity for retraining.
The remainder of this paper is structured as follows. Section 2 reviews past research in the field of web tracker detection. Section 3 introduced definitions, our research questions, and the analyzed data sets. Section 4 explores HTTP/S response headers from Chrome, Firefox, and Brave data sets, providing general findings and identifies potential challenges for the ML pipeline. Section 5 describes the design and implementation of the ML pipeline, including selected models, metrics, and baselines. Section 6 presents the evaluation of the trained classifiers. The penultimate Section 7 discusses our results and limitations. Section 8 concludes the paper.
2. Related Work
This section provides an overview of existing research on web tracker detection based on machine learning. To highlight how our proposed classifier differs from prior work, we focus this overview on classifiers that rely solely on HTTP protocol information. For the sake of completeness, however, we summarize other approaches at the end of the section.
2.1. Classification Based on HTTP Information
Approaches to the detection web trackers with the support of machine learning date back to the early 2000s. In 2010, Yamada et al. (Yamada et al., 2010) were the first to use temporal link analysis to build a graph based on hosts related to HTTP requests from the traffic generated by corporate network. Their classifier considers temporal information such as the visiting time of a specific website or host. Yamada et al. consider this a discriminating criterion based on the assumption that users spend less time on websites of web trackers, as trackers generally appear as third parties that are contacted briefly in the context of other websites rather than first parties that users actively visit.
In 2013, Bau et al. (Bau et al., 2013) described and discussed the usefulness of machine learning as an effective counter to web tracking. They outlined the problems with manually curated filter lists - problems that still hold true today, such as the reliance on manual curation and the inflexibility of static lists. Beyond building an Elastic Net model focused on traces of content extraction in the Document Object Model and HTTP headers to show the advantages of automated approaches, Bau et al. laid down requirements for future machine learning models as well as problems facing researchers in this domain.
A year later, Bhagavatula et al. (Bhagavatula et al., 2014) tried to identify ad-related URLs, a topic closely connected to the web tracking detection problem. To this end, they crawled the Alexa top 500 websites to compile an extensive dataset with six distinct feature sets based on the URLs themselves as well as additional metadata. Their analysis approach is based predominantly on Support Vector Machines (SVMs) with different kernels, achieving average accuracy and precision scores above 89%. Like Bau et al. before them, Bhagavatula et al. recognized the need for machine learning models to replace filter lists and put forward additional requirements for production-ready machine learning approaches, such as optimized training times.
In 2015, Gugelmann et al. (Gugelmann et al., 2015) proposed an alternative approach by deriving several statistical features from HTTP traffic, e.g. up- and download volume. In contrast to prior studies, they thoroughly described their model-building process and emphasized their feature engineering. Their feature selection process included the use of Pearson’s correlation coefficient as well as the consideration of a feature’s entropy. In addition, they employed an unspecified classification tree, among other classifiers, to find an optimal Naïve Bayes model that achieved a precision of 83% and was able to correctly identify previously unseen trackers.
In the same year, Li et al. (Li et al., 2015) proposed TrackAdvisor, a system to classify HTTP requests based on the collective statistics of cookies contained therein. Using FourthParty, they captured 563,031 requests and identified 22,270 cookies that were sent to third parties. Rather than relying solely on filter lists, they used a multi-step process with an emphasis on manual inspection to label the captured data. Notably, the training and test sets created from the captured data comprised only 500 requests each and had a balanced class distribution, which contrasts with the distribution found in the original data. Nevertheless, Li et al. used these sets to train a SVM that ultimately achieved a precision of 99.4% and a recall of 100%. Like Gugelmann et al., they point out the usefulness of the HTTP Referrer header for third-party tracker detection.
In 2016, Dudykevych and Nechypor (Dudykevych and Nechypor, 2016) also used cookies as the basis for their classifiers. In contrast to Li et al., however, they developed their own crawler, had a larger data set of 550 000 cookies, and developed a different feature set. The latter is only described briefly, but information from the cookies themselves and the HTTP requests were used as features. It can be assumed that the features are numerical and categorical, as the authors used Principal Component Analysis and correlation, which require numerical data, as well as binary encoding of categorical features. The resulting classification accuracies ranged from 82.01% for Naïve Bayes to 95.43% for the regularized logistic regression method.
The analysis of HTTP traffic for and with machine learning models is also an important topic in web security research, albeit with a different target (i.e. label) than web privacy research. In a 2019 study, McGahagan et al. (McGahagan et al., 2019) analyzed response headers from 45 874 websites and extracted 672 features, resulting in a final feature vector with 22 properties. In addition, they tested whether over- and undersampling had any effects on the model’s performance. The coverage of metrics, models, different training and tuning combinations, and extensive feature engineering documentation can be considered standard among machine-learning practitioners but is somewhat rare among prior related research. An important contribution of this paper is the in-depth presentation of the usefulness of HTTP response headers for automated ML models.
Unlike most papers discussed in this section, which have used data from the Firefox browser to test the performance of supervised learning models, Din et al. (Din et al., 2020) used Deep learning (DL) techniques for Brave and other Chromium-based browsers. Although their approach is geared towards the image-based detection of advertising in websites rather than web tracking, their incorporation of multiple browsers is highly relevant for models that classify web trackers.
Cozza et al. (Cozza et al., 2020) developed a system that not only classifies HTTP requests but also automatically updates the filter list. Although this study is not the first to detect previously undocumented trackers using machine learning techniques, the deployment of a tracker-detecting browser extension that combines a filter list with a machine learning models and updates its ground truth marks an important step forward in tracker detection technology.
Guarino et al. (Guarino. et al., 2020) built upon prior research by Gugelmann et al. and Li et al. by training multiple model based on combinations of features used by these studies, including different HTTP headers and the ratio of upstream and downstream data transmissions. Their top performer is a Random Forest model with an average accuracy of 92%. It should be noted, however, that this study is based on a relatively small dataset of just 1000 observations. Nevertheless, the authors astutely emphasize that content delivery networks (CDNs) are a primary cause of false positives in the domain of web tracker detection due to similarities in traffic patterns between CDNs and third-party trackers.
Taking advantage of the observation that HTTP-based communication between clients and servers often comprises cascades of multiple HTTP messages, Iqbal et al. (Iqbal et al., 2022) conducted an extensive study that uses entire request chains rather than individual requests as the basis for classifying web tracking. Notably, they were among the first to consider responses as well as requests by developing three distinct feature sets: the first is based on the entire request chain, the second relies on information pertaining to individual requests and the third considers response headers and metadata. By training a Random Forest model on five distinct datasets with different configurations from separate crawls, Iqbal et al. achieved accuracies above 98%. However, if the model was trained on only one crawl and tested on the others, the accuracy dropped by up to 14%, highlighting the difficulties of applying models to datasets with varying configurations, e.g. data collected across different browsers.
2.2. Other Detection Approaches
Beyond the HTTP-focused approaches outlined above, a variety of classification methods utilizing different data sources for the extraction of distinctive criteria have been proposed and evaluated. These include, classifiers focused on JavaScript code and API access events that aim to identify device fingerprinting techniques, pure URL-based classifiers, CNAME cloaking detectors, and graph-based classifiers that leverage the topology of the Web to identify third-party tracking in particular. In the following paragraphs, we will briefly highlight notable research on these alternative approaches.
Focusing on JavaScript API calls, Wu et al. (Wu et al., 2016) developed DMTTrackerDetector, a system using decision trees in which observed API calls serve as features. The system, comprising a 717-dimensionial feature vector, achieved 97.8% accuracy and addressed obfuscation issues, showing the distinct JavaScript API usage by trackers and non-trackers.
Bahrami et al. (Bahrami et al., 2021) refined the scope of their analysis by focusing on device fingerprinting rather than API calls in general. Their approach combines graph analysis with supervised and unsupervised learning to detect web trackers in historical JavaScript file data collected between 2010 and 2019 by creating a graph that includes temporal aspects and the output of AST-based keyword extraction. Their method, which includes nine graph-related features and graph embedding, as well as a clustering model, yielded accuracies up to 89.13% and identified distinct clusters for fingerprinting and functional APIs, highlighting the viability of such an approach.
Kalavri et al. (Kalavri et al., 2016) also proposed a graph-based approach, using neighborhood analysis and label propagation on a bipartite graph representing connections between first-party websites and third-party services to detect web trackers, achieving over 97% accuracy. They analyzed six months of data, finding that most trackers are closely linked, with precision scores ranging from 64 to 92%.
Combining these graph-based approaches, Castell-Uroz et al. (Castell-Uroz et al., 2021) proposed a tripartite network graph to identify third-party tracking resources and their origin domains within first-party websites. By applying this approach to a dataset comprising over 90 thousand domains, they were able to detected new trackers through their hash value popularity and dirt level, finding that higher values correlated with the likelihood of a given resource to be tracking-related.
Conversely, Metwalley et al. (Metwalley et al., 2015) developed an unsupervised algorithm analyzing URL queries and HTTP request headers for identifying trackers. Their method, applied to 200 Alexa-ranked websites, identified 34 new trackers and successfully supported the analysis of cookie-matching practices, though it had limitations in detecting all known trackers.
Finally, Yu et al. (Yu et al., 2016) proposed a browser extension using k-Anonymity principles to classify third-party trackers based on data sent to users. Leveraging data from 200,000 German users of the Cliqz browser, their system outperformed traditional blockers like Disconnect, which had a 38.8% false positive rate.
2.3. Summary
Despite over a decade of dedicated research, web tracking is as prevalent as ever, new tracking methods are developed, and the most widespread form of protection, namely blockers and their filter lists, exhibit several key problems. Filter lists, which are also used as the ground truth for labeling in many studies related to web privacy, are often maintained by a small group of users, contain irrelevant rules, are rather slow to add new rules, and can simply be circumvented by changing URLs. In response, researchers have developed a wealth of automated approaches, utilizing HTTP messages, JavaScript, or the HTML files of websites themselves to classify third parties as trackers.
A crucial component of this ongoing research effort is the availability and quality of data on which to train and evaluate machine learning approaches.
Despite this, HTTP response headers are often neglected in this regard. There are cases where authors document the collection of HTTP responses (Acar et al., 2013) but not their actual use. In other cases, the data is collected, but only used in support of other approaches such as the classification of cookies (Lerner et al., 2016; Englehardt et al., 2015).
Additionally, the majority of studies are based on datasets collected using just one browser, most commonly Firefox. Cassel et al. (Cassel et al., 2022) discuss the need for cross-browser examinations to avoid biased datasets. This is supported by research that shows that characteristics such as JavaScript API calls differ between browsers (Yang and Yue, 2020).
In summary, our review of related research has shown that, although the application of machine learning to the field of web tracking identification and classification has been investigated in numerous studies, a research gap still exists with regard to cross-browser performance of classifiers and the usefulness of HTTP response data for classification purposes.
3. Approach
In this study, we address the research gap identified in our review of related research by evaluating the (cross-browser) performance of commonly used machine learning classifiers that are trained on HTTP/S response headers. In this section, we discuss our underlying research questions and present the approach we used to address them.
3.1. Preliminaries
A core principle of automated web tracker detection is the definition of what constitutes a tracker. The focus hereby lies on the tracking entities and not on the activity of tracking. Our definition of what constitutes a tracker or non-tracker is coupled to our ground truth to label each dataset. We define them as follows:
[…]
$\text{match}(h,r)$ :$:\begin{cases}1&\text{if }h\text{ matches rule }r\\ 0&\text{otherwise}\end{cases}$
$T$ =$=\{h\in H\mid\exists r\in R:\text{match}(h,r)=1\}$
$NT$ =$=\{h\in H\mid\forall r\in R:\text{match}(h,r)=0\}$
In other words: if the remote hostname (i.e. the sender) of a response matches at least one rule in our aggregate list compiled from the EasyList and EasyPrivacy filter lists, we label the response as a tracker.
For the sake of simplicity, we will use the terms header and response header interchangeably to refer to HTTP/S response headers throughout the remainder of the paper, unless explicitly stated otherwise.
3.2. Research Questions
The main goal of this study is to analyze whether response headers are useful for automated web tracker detection. Based on the ML approaches of prior research and the observed areas to further extend our empirical understanding and detection of web trackers, we formulate the following research questions that guide us throughout this study:
  1. (1)
    What are the differences and similarities between web trackers and non-trackers with regard to their response headers?
  2. (2)
    Can trained classifiers detect web trackers with only response header features while achieving an F1-Score higher than 0.9 in an imbalanced setting?
  3. (3)
    How large is a classifier’s performance degradation when it is trained on one browser’s HTTP traffic data and then applied to other browsers?
To answer these research questions, we develop a semi-automated ML pipeline that transforms HTTP traffic datasets collected with T.EX (Raschke et al., 2019) and uses it to train several supervised binary classifiers. We use reliable data that was compiled by prior research to facilitate future comparisons between new classifiers.
To test the robustness and reliability of our classifiers, we train them on HTTP data from Chrome and apply them to different datasets with a variety of configurations, including other browsers (latitudinal comparison) as well as different time frames within the same browser (longitudinal comparison). Furthermore, past research observed an uneven distribution of trackers and non-trackers, which is why we train our classifiers with an imbalanced distribution, to assess how effective they perform with the likely true distribution.
$\text{Chrome}_{22}$ $\text{Firefox}_{22}$ $\text{Brave}_{22}$ $\text{Chrome}_{23}$
T NT T NT T NT T NT
#Responses 256 202 602 516 309 391 526 865 2969 561 694 206 747 596 066
#Responses in % $\approx 0.3$ $\approx 0.7$ $\approx 0.37$ $\approx 0.63$ $\approx 0.5$ $\approx 99.5$ $\approx 0.26$ $\approx 0.74$
#Unique headers in total 2213 7577 2516 7322 237 7348 2164 7398
Unique Headers per response ($Q1,\tilde{\mu},Q3$) 10,13,16 12,15,19 12,15,18 13,16,19 13,18,20 12,15,19 10,13,16 13,15,19
Table 1. Overview of the collected data sets from (zen, 2023), split by data sets as well as tracker (T) and non-trackers (NT).
Figure 1. This figure illustrates the usage headers across four different browsers. The left plot (a) displays the frequency distribution of headers for each browser, showing how often each header is present in the dataset. The right plot (b) presents a boxen plot of the same data, which provides a more detailed view of the distribution, revealing variations and trends among the browsers.
3.3. Data Collection
We used existing data sets, containing labeled HTTP response headers by EasyList and EasyPrivacy. Following suggested good practices by (Demir et al., 2022), the crawl should fulfill three citeria: (i) repeated crawls across multiple browsers, namely Chrome, Firefox, and Brave, (ii) the documentation of the geographical region as well as the technical environment, and (iii) web sites from the Tranco ranking as it proved to be less biased than the Alexa ranking (Pochat et al., 2018). A previously crawled data set by Raschke using T.EX - The Transparency EXtension, fulfills all three criteria and is described as follows:
Multiple simultaneous and stateful crawls (6 x Chrome, 6 x Brave, 6 x Firefox) of the Tranco top 10K websites (as of 16th of May 2022) were performed on the 12th of August 2022 with T.EX […]. Measurements had been carried out on 18 Amazon Web Services instances (c5.large) running Windows Server. All instances were launched in Frankfurt, Germany (eu-central-1).
After completing the crawls, each instance’s extension storage was extracted […]. Note: 2 of the 6 Firefox instances crashed during the crawl. Therefore, only 4 Firefox datasets are available (zen, 2023).
To assess a potential concept drift and therefore model degradation over time, we crawled one more data set from 2023 for Chrome. We followed a similar methodology as described in (zen, 2023), using T.EX with Chrome to explore the top 10K websites from the Tranco ranking as of March 29th, 2023.
Our reasons to choose only Chrome for a new data collection are as follows: Chrome is the most widely used in the real world with a market share of 66.13% on desktops, whereas Firefox has 7.1% (sta, 2023) and Brave has less than 1% (kin, 2023). Based on this distribution, the Chrome data set serves as the foundation for training to reflect and reach the majority, i.e. using data that most users encounter themselves. In addition, web privacy research often relies on a simulated Firefox browser with Selenium due to the use of OpenWPM, leaving space for other browsers of interest. In addition, it helps us to answer the question whether a model based on Chrome data can be applied to Firefox or Brave while achieving high performance metrics?.
4. Initial Response Header Analysis
To get a better understanding of the data underlying our analysis before designing our machine learning pipeline, we conduct an initial analysis of the data, the results of which we present in this section. We first explore response headers across all three browsers to find empirical evidence of their potential relation to trackers. Furthermore, we identify the characteristics of these datasets to apply the appropriate methods in our ML pipeline. We use these insights to develop our classifiers to evaluate the effectiveness of response headers for web tracker detection.
4.1. Descriptive Analysis of Headers
A descriptive analysis is useful to gain an understanding of the data structure and the distribution of classes and features within the dataset. Table 1 summarizes descriptive metrics for each dataset. We start with the datasets $\text{Chrome}_{22}$, $\text{Chrome}_{23}$, and $\text{Firefox}_{22}$. All three datasets show similar results for the distribution between trackers and non-trackers, the number of unique headers in total and per response. For example, we can observe a small quantitative change in the number of unique headers between $\text{Chrome}_{22}$ and $\text{Chrome}_{23}$. $\text{Chrome}_{22}$ lists a total of 7577 unique headers, while $\text{Chrome}_{23}$ has a slightly lower total of 7398 for non-trackers. When comparing trackers, $\text{Chrome}_{22}$ had 2213 unique headers, which is nearly the same as $\text{Chrome}_{23}$’s count of 2164. On average, non-trackers have more unique headers as well as a higher median $\tilde{\mu}$ and interquartile range than trackers.
The $\text{Brave}_{22}$ browser with its standard configuration is the only one that has a low number of trackers and unique headers in tracking-related response messages. Interestingly, the $\tilde{\mu}$ equals 18 for trackers and 15 for non-trackers, which represents the only instance where trackers had a higher number. Although the dataset has almost only non-trackers, the number of unique headers is similar to the other datasets.
Figure 1 shows a frequency plot, which is a distribution graph that shows how often each header appears in the dataset. It uses a logarithmic scale to manage the wide range of frequencies, which is a common approach when data spans several orders of magnitude. This plot typically demonstrates a power-law distribution, characterized by a steep drop-off where a small number of headers are occurring frequently, while the majority are rare. This type of distribution is common in HTTP data, where few headers dominate, and many others have a long tail of low frequency. This sparseness indicates that while there are many different headers, only a few are consistently used across different websites and by different browsers.
The boxen plot graph of Figure 1 offers a visual representation of the distribution of header frequencies across different browsers and versions. The spread and median values suggest that the distribution of headers is not uniform across browsers, with some browsers having a higher tendency to use certain headers than others. Furthermore, if the median line in the boxen plot is closer to the bottom of the box, it suggests a right-skewed distribution, aligning with what the frequency plot shows — a few headers with high frequency and many with low frequency.
Figure 2. The Venn diagram reveals that a core set of 5672 headers are common across Chrome, Firefox, and Brave browsers across two years. Unique headers are present in each browser, with $\text{Chrome}_{23}$ having the highest number of unique headers, which may reflect browser-specific features or experimental headers. Firefox had a higher header similarity with Chrome than with Brave.
Figure 2 represents a Venn diagram and provides a direct comparison of response headers between different browser versions, highlighting data and feature similarity. The intersection numbers, such as the 5672 headers common to all browsers, suggest a significant overlap in features across different platforms. However, the presence of headers unique to each browser or shared by only two browsers (e.g., 1427 unique to $\text{Chrome}_{22}$ and 154 unique to $\text{Brave}_{22}$) indicates that while there is a core set of features, there is also a notable amount of variation which can be critical for web tracker detection.
The Empirical Cumulative Distribution Function plot for $\text{Chrome}_{22}$ and $\text{Chrome}_{23}$ in Figure 3 row (A) show that the Content-Length values for non-trackers are generally higher than those for trackers, which suggests that non-trackers tend to send larger responses. This is consistent across both browser versions, indicating that while browsers evolve, certain characteristics like the smaller size of tracker responses remain consistent. The second row (B) plots the feature similarity for the X-XSS-Protection header, showing that header values can be very similar between the datasets, i.e., the training and test sets. It further supports our observation that the datasets are similar and that trackers and non-trackers have differences across the usage of HTTP header fields and values. Although we present only few selected examples here, they should highlight the general challenges that come with response messages datasets.
Figure 3. Row (A) presents the ECDFs for both Chrome datasets for the Content-Length header values. We set a cut-off value at 10 000 to highlight the core observation and more than half of the responses have a value are below this threshold – $\tilde{\mu}_{T22}=62$, $\tilde{\mu}_{NT22}=8068$ and $\tilde{\mu}_{T23}=45$, $\tilde{\mu}_{NT23}=8021.5$. Row (B) shows how similar the values across Chrome, Firefox, and Brave are for the X-XSS-Protection header.
Additional characteristics of the datasets include the high amount of missing values per header. One explanation might be that servers do not need to use every header field that exists, and it is common practice to use developer headers that are not part of the standard, i.e. they are not registered in the Request for Comments 4229111RFC 4229, https://www.ietf.org/rfc/rfc4229.txt, accessed: November 30, 2023. There are multiple headers that contain mostly unique values, e.g., Date. Lastly, according to the used filter lists, most HTTP response messages are not related to any tracking activities. This entails a class imbalance to be handled carefully to derive meaningful results from the final models.
5. Methodology
Based on our findings from the data exploration in Section 4, we propose our methodology to assess the effectiveness of classifiers based on response headers.
5.1. Problem Formulation
We consider the problem of binary classification, where we aim to predict the presence of web trackers based on response headers. The headers, represented by the feature vector $x\in\mathbb{R}^{d}$, are nominal and can be both discrete and continuous variables. These variables are derived from HTTP headers of four different browser types, denoted as the set $B=\{\text{Chrome},\text{Firefox},\text{Brave}\}$. For this study, we specifically consider versions $\text{Chrome}_{22},\text{Chrome}_{23},\allowbreak\text{Firefox}_{22},\text{Brave}_{22}$.
Given a training set $\mathcal{D}_{\text{train}}=\{(x_{1}y_{1}),\ldots,(x_{n},y_{n})\}$, where $x_{i}$ is a feature vector of the $i$-th HTTP header and $y_{i}\in\{0,1\}$ represents the associated class label (with ’1’ indicating the presence of a web tracker), our goal is to train a set of classifiers $f_{j}:\mathbb{R}^{d}\rightarrow\{0,1\}$ for $j\in M$ that minimizes the empirical risk:
(1) $R(f_{i})=\frac{1}{N}\sum_{i=1}^{N}L(y_{i},f_{i}(x_{i})),\quad j\in\{1,\ldots,M\},$
where $L(y_{i},f_{i}(x_{i}))$ is the loss incurred by predicting $f_{i}(x_{i})$ when the true label is $y_{i}$. The specific form of the loss function $L(\cdot)$ will be dependent on the particular classifier $f_{j}$ used.
We consider datasets represented by $D_{s}$, where $s$ can pertain to training, testing, validation, or calibration subsets of the data. We observe several challenges from our data exploration:
  1. (1)
    High Cardinality: The feature vectors $x_{i}$ in our datasets $D_{s}$ can exhibit high cardinality for certain features. This problem can pose challenges in terms of model training and encoding.
  2. (2)
    Sparseness: The feature vectors $x_{i}$ in our datasets $D_{s}$ are represented as sparse vectors, with each vector defined as:
    $x_{i}=\{v_{1},v_{2},\ldots,v_{d}\}$
    Here, a significant proportion of the entries $v_{j}$ within $x_{i}$ are either zero or represent missing data, leading to the overall sparseness of $D_{s}$. Such sparseness can introduce computational complexities and may challenge the performance of certain machine learning algorithms not optimized for sparse data.
  3. (3)
    Class Imbalance: For datasets collected from a given browser $B$ and represented as $D_{s}$, the class distribution can exhibit imbalance. The majority class is denoted by $\alpha_{B}$, and the minority class is represented by $\beta_{B}$, satisfying $\alpha_{B}+\beta_{B}=1$. The degree of this imbalance can vary based on the browser as shown in Table 1, with some showing a more pronounced disparity between $\alpha_{B}\text{ and }\beta_{B}$. Such imbalance can introduce bias to models, favoring the majority class and potentially undermining the accurate prediction of the minority class.
  4. (4)
    High Dimensionality: With a high dimensionality $d$ and a high number of observations $n$ per dataset, the classification is susceptible to the Curse of Dimensionality, potentially leading to overfitting and necessitating intricate dimensionality reduction techniques. We have to apply a dimensionality reduction process that achieves a reduced feature space $X_{i}^{\prime}\subseteq X$, entailing a reasonable training time.
  5. (5)
    Data Similarity: The datasets $\mathcal{D}_{s}$ collected across different browsers can exhibit pronounced similarities in their feature distributions and values. Given the high dimensionality of our feature space, exhaustive pairwise comparisons can be computationally expensive and time-consuming, highlighting the challenge of accurately gauging data similarities in such a complex space.
5.2. Machine Learning Pipeline
Our machine learning pipeline, which we present in the following, addresses the previously stated problems for the classification of web trackers. Many different approaches have been applied in order to uncover web trackers. Another closely related field to web privacy is web security research. Although they seek to detect different entities, both have the exploitation of HTTP data in common. Typically, a dataset goes through extensive data processing and feature engineering to find useful features, which includes a thorough analysis of the values within each feature. Despite this common conception, Laughter et al. (Laughter et al., 2021) were interested if features that merely represented the presence of an HTTP header in a binary fashion were usable in a classifier to detect malicious HTTP requests based on HTTP headers and URLs. In the end, some of their classifiers performed well with accuracies above 90%. Similar results were observed in (Bhagavatula et al., 2014) and partly by (Dudykevych and Nechypor, 2016).
This method is interesting on two levels: (i) it cuts down the time of feature engineering, which is fairly complex in this use case as HTTP headers are mainly categorical data, and (ii) it should help when the models are deployed and used in production, e.g. if HTTP data is received, it should be evaluated in real time and checking the presence of headers is faster than elaborate processing before evaluation. Therefore, the binary representation of HTTP headers as a feature will be tested.
First and foremost, we train our classifiers based on the $\text{Chrome}_{22}$ dataset and use all other datasets as test sets, which is in accordance with RQ3. Generally, each step in our pipeline focuses on reducing the dimensions of the datasets, as most headers are rarely used, thus reducing potential overfitting. This approach further reduces the sparseness and potential features with high cardinality. The latter challenge is solved by employing the binary-encoding method as it flattens the respective feature to two possible values.
The sparseness and high dimensionality are related to each other, in a sense that most headers and their corresponding values are rare or even unique. We apply multiple methods to reduce the dimensionality: (i) a low variance filter to remove headers with only one value, (ii) the removal of headers with a missing value ratio of one, (iii) headers related to only one label are dropped, and (iv) Approximate string matching, sometimes referred to as Fuzzy Matching to combine header fields with a similarity of at least 80% based on a weighted ratio of multiple distance metrics, e.g., Damerau Levenshtein or Hamming. All of this reduces the sparseness further and binary-encoding simplifies the sparse matrix.
Although a multitude of different resampling techniques exists to combat class imbalance, the question of whether this step is actually necessary remains. Some disadvantages of such methods are, for example, the removal of observations, the replication of existing observations, or the creation of synthetic samples. Selecting and applying a sampling method can lead to overfitting and overoptimism as well (Santos et al., 2018), notwithstanding the question of whether it is even feasible to train the models on a distribution that was not observed in real-world data. Our pipeline preserves the observed distribution as the imbalance is mild (goo, 2023) in the $\text{Chrome}_{\textbf{22}}$ dataset, i.e., a sufficient number of samples for each class exists, which allows proper training to differentiate between classes. Furthermore, we compute metrics that consider or are not affected by the imbalance of the datasets. The same applies for most of the selected ensemble models.
Fuzzy matching is especially useful due to the presence of hundreds of similar headers within our data. During our initial data analysis deiscussed in Section 4, we identified many examples where headers with similar purpose were written or formatted differently, e.g. content-lenght, cteonnt-length, or ntcoent-length, instead of content-length. Accordingly, simply comparing the header names is not enough, which is why we automatically assess the value similarity to reduce potential mismatches.
Lastly, we apply Isotonic Regression instead of Platt Scaling to calibrate the probabilities outputted by our models, following the results from (Niculescu-Mizil and Caruana, 2005). This is done by fitting an Isotonic Regression model to the models’ predictions and the true labels, and using this to adjust the predicted probabilities. The calibration function for model $i$ is denoted as $C_{i}(f_{i}(x))$.
We summarise our pipeline in Figure 4, including the data collection. Note that we did not apply any data imputation to keep the final datasets as close to the original collected datasets as possible. A validation set for hyperparameter tuning was not created because we want to assess the effect of calibration on its own. Moreover, tuning requires substantially more computational power, e.g. by using Grid Search.
Figure 4. The parameterized pipeline automatically processes the collected datasets and trains the classifiers.
5.3. Classification Models
As discussed in Section 2.1, prior studies (Cozza et al., 2020; Li et al., 2015; Bhagavatula et al., 2014; Gugelmann et al., 2015) often used similar classifiers despite the broad range of available models. We build upon these established selections, refining our choice of models based on three conditions: (i) the models should be trainable in a reasonable amount of time, (ii) they should be applicable to this specific use case, and (iii) they should preferably be ensemble methods, although we include a handful of non-ensemble models that were used in past research to allow for direct comparisons between our results and those of related studies. We prefer ensemble methods as they usually perform better than their counterparts (Dietterich, 2000), leading to their popularity over the past few years. Nevertheless, to our best knowledge, they have not been widely used in past web privacy research, except for Random Forests (RF). In Table 3, we provide an overview of our selected models and their use in related work.
Model Ensemble Past Research
Decision Tree No (Gugelmann et al., 2015; Dudykevych and Nechypor, 2016; Raschke et al., 2023)
Random Forest Yes (Cozza et al., 2020; Wu et al., 2016; Guarino. et al., 2020; Iqbal et al., 2022; Raschke et al., 2023), (Dudykevych and Nechypor, 2016) (U), (McGahagan et al., 2019)
Extra Trees Classifier No Not used
Logistic Regression No (Bhagavatula et al., 2014; Gugelmann et al., 2015; Dudykevych and Nechypor, 2016; Wu et al., 2016; Raschke et al., 2023), (McGahagan et al., 2019)
Naïve Bayes (Gaussian or Bernoulli) No (Yamada et al., 2010; Bhagavatula et al., 2014; Gugelmann et al., 2015; Dudykevych and Nechypor, 2016; Wu et al., 2016; Guarino. et al., 2020)
Gradient Boosting Yes (McGahagan et al., 2019), (Raschke et al., 2023)
Light-Gradient Boosting Machine Yes Not used
Adaptive Boosting Yes (Raschke et al., 2023)
Hist-Gradient Boosting Yes Not used
eXtreme Gradient Boosting Yes (Raschke et al., 2023)
Table 2. Overview of the selected machine learning models and the number of times they were used in past research from Section 2.1. One paper did not exactly specified which ensemble model was used, thus, the most likely model was marked with Unspecified (U). Papers from web security are underlined.
Other supervised learning methods, e.g., Multi-layer Perceptrons (Cozza et al., 2020; Guarino. et al., 2020), KNN, and especially SVM, could have been used. The latter was successfully applied in multiple papers (Cozza et al., 2020; Li et al., 2015; Bhagavatula et al., 2014; Guarino. et al., 2020). However, both suffer in terms of time and computational complexity, e.g. SVM has a train time complexity of $O(n^{2})$ where $n$ is the number of observations (Steinwart and Christmann, 2008). In addition, deep-learning models are not considered as the focus is on supervised classifiers and simple rather than complex models. A graph-based approach, used by several papers (Kalavri et al., 2016; Castell-Uroz et al., 2021) will not be pursued because it does not fit well with the targeted approach, and direct connections between individual HTTP responses will not be analyzed, e.g. through grouping based on their URL or building a graph-based on their order.
5.4. Baseline Models
We compare our classifiers against three baseline models and multiple metrics that we define in Section 5.5. We select one tree classifier, the Decision Tree (DT), that is also the building block of six selected classification models. Two simplistic and relatively robust baselines, namely Naïve Bayes (NB) and Logistic Regression (LR).
Additionally, it is good practice to compare against a baseline approach that differs from the presented classifiers while using the same dataset. Raschke et al. (Raschke et al., 2023) presented a graph-based approach named T.EX-Graph where edge and node attributes as well as centrality metrics served as features. Furthermore, the authors included information from the responses but did not pursue classifiers that are solely based on response headers. This comparability should help in assessing the performance gains and the relevance of response header values for automated web tracker detection. However, the authors applied Synthetic Minority Oversampling to balance their dataset. We focus on their results using the FQDN dataset, where hosts are modeled as fully qualified domain names (FQDN) because the tracker distribution is similar to ours with a NT to T ratio of 68:32.
5.5. Evaluation Metrics
We measure the performance of our classifiers according to nine different metrics. Several metrics are common in the web tracking literature, e.g., accuracy, precision, or F1-Score, while others were selected due to their applicability to our classification problem and to understand the relationship with the underlying data in more detail. Additional metrics are reported in the Appendix, e.g., the confusion matrix because its interpretation is already captured by other metrics more concisely. Moreover, we calculate confidence intervals (CI) for these metrics using bootstrap sampling with 599 samples (Wilcox, 2010).
Metric Past Research
(Avg.) Accuracy (Gugelmann et al., 2015; Dudykevych and Nechypor, 2016; Raschke et al., 2023; Cozza et al., 2020; Guarino. et al., 2020; Iqbal et al., 2022; Bhagavatula et al., 2014), (McGahagan et al., 2019)
Precision (Gugelmann et al., 2015; Raschke et al., 2023; Iqbal et al., 2022; Bhagavatula et al., 2014; Li et al., 2015)
Recall (Gugelmann et al., 2015; Raschke et al., 2023; Iqbal et al., 2022; Li et al., 2015)
F1-Score (Gugelmann et al., 2015), (Raschke et al., 2023)
AUC-ROC (Gugelmann et al., 2015), (Bhagavatula et al., 2014), (McGahagan et al., 2019)
Confusion Matrix (Guarino. et al., 2020)
MCC (McGahagan et al., 2019)
AUCPRC Not used
Log-Loss Not used
Table 3. Overview of performance metrics and the number of times they were used in past research from Section 2.1.
The first metrics that are commonly used in web privacy research are Accuracy, Precision, Recall, False Positives (FP), True Positives (TP), False Negatives (FN), and True Negatives (TN). Two accuracy scores can be calculated, one from the perspective of trackers, the positive label, and one from non-trackers, the negative label. However, accuracy is not particularly useful with imbalanced datasets because a model could always predict the majority class and achieve accuracy with the same value as the distribution of said class. We calculated additional, more reliable metrics to quantify the performance of a classifier with an imbalanced dataset as follows:
Balanced Accuracy:
Broderson et al. (Brodersen et al., 2010) introduced the Balanced Accuracy (BACC) metric for imbalanced datasets in 2010 as an alternative to average accuracy scores, which are, according to the authors, problematic for confidence intervals (CI) and imbalanced datasets. The metric comprises recall and TNR and is defined as:
(2) $BalancedAccuracy=\frac{1}{2}\left(\frac{TP}{TP+FN}+\frac{TN}{TN+FP}\right)$
F1-Score:
Another metric that is commonly used for classification models is the F1-Score which expresses a model’s performance in one value, despite the ambivalent nature of precision and recall. Compared with BACC, it emphasizes positives more, whereas BACC takes negatives more into account, which can result in two very different results. The F1-Score is mathematically defined as the harmonic mean between recall and precision:
(3) $F_{1}\text{-}Score=\frac{2\cdot Precision\cdot Recall}{Precision+Recall}$
Log-loss Score:
This metric, sometimes referred to as cross-entropy-loss, is particularly useful for classification problems as it focuses on the predictive performance of a model. The function is applied to each observation individually, and then the average over all elements is calculated. It describes how large the error between a prediction probability and the true label is, and a value closer to zero represents a very good score. However, this metric can be influenced by the imbalance in a dataset, i.e. the baseline log-loss gets smaller with an increase in imbalance, and the log-loss metric should always be compared to its baseline value, e.g. the baseline log-loss for a dataset with a 70:30 ratio equals 0.611, and 0.325 for datasets with a 90:10 ratio (tds, 2023). All in all, the model’s log-loss should be lower than its baseline. The math formula is defined as:
(4) $L_{log}(y,p)=-[y\cdot log(p)+(1-y)\cdot log(1-p)]$
where $y$ stands for the label and $p$ for probability.
Matthews-Correlation-Coefficient:
One metric is often not enough to fully comprehend the performance of a model and if it is even suited for binary classification problems. Matthews-Correlation-Coefficient (MCC) is becoming more popular in recent years, although it was introduced in 1975 (Matthews, 1975) – it is defined as:
(5) $MCC=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)\cdot(TP+FN)\cdot(TN+FP)\cdot(TN+FN)}}$
However, it is rarely used in web privacy research compared to other fields like web security research. MCC ranges from -1 to +1 and can be understood as a correlation coefficient between the true and predicted label. One particular advantage is the inclusion of values from the confusion matrix, i.e. the MCC will only achieve a high value if both class labels were predicted with high accuracy – which in turn presents the MCC as a meaningful metric for imbalanced datasets (Boughorbel et al., 2017). This was further substantiated by Chicco et al. (Chicco and Jurman, 2020) in 2020, and the authors outlined how accuracy and F1-Score were not reliable measures for imbalanced datasets.
ROC-AUC:
Our next metric is the Receiver Operating Characteristics (ROC) curve, complemented by the Area Under the curve (AUC) value. The former describes a classifier’s performance at various thresholds by plotting the TPR on the ordinate and FPR on the abscissa. The latter metric expresses how well the classifier is able to separate between both class labels.
AUPRC:
The last metric in our study is the Area Under the Precision-Recall Curve (AUPRC). This metric is particularly relevant in the context of a supervised binary classification problem, especially when dealing with imbalanced datasets. The Precision-Recall curve is a plot that illustrates the trade-off between the precision (positive predictive value) and the recall (true positive rate) for a classifier at various thresholds. Precision is defined as the ratio of true positives to the sum of true and false positives, while recall is the ratio of true positives to the sum of true positives and false negatives. The AUPRC is an aggregated measure of this curve, representing the classifier’s ability to correctly identify positive instances across different thresholds. In imbalanced settings, where one class is significantly underrepresented, the AUPRC is a more informative and reliable metric than ROC-AUC, as it focuses on the performance of the classifier on the minority class, which is often of greater interest (Saito and Rehmsmeier, 2015).
6. Experiments
Model Accuracy Log-Loss ROC-AUC AUPRC BACC F1-Score Precision Recall MCC
LR 0.892 0.252 0.86 0.902 0.86 0.812 0.847 0.78 0.738
[0.891;0.894] [0.25;0.255] [0.858;0.862] [0.9;0.904] [0.858;0.862] [0.81;0.815] [0.843;0.85] [0.777;0.784] [0.735;0.742]
GNB 0.853 2.47 0.832 0.792 0.832 0.759 0.74 0.779 0.654
[0.851;0.854] [2.436;2.514] [0.829;0.833] [0.788;0.797] [0.829;0.833] [0.756;0.762] [0.737;0.745] [0.776;0.782] [0.65;0.657]
BNB 0.851 0.617 0.833 0.849 0.833 0.759 0.732 0.789 0.652
[0.849;0.852] [0.608;0.627] [0.83;0.835] [0.847;0.852] [0.83;0.835] [0.757;0.763] [0.729;0.736] [0.784;0.793] [0.648;0.657]
DT 0.956 0.546 0.947 0.943 0.947 0.926 0.928 0.925 0.895
[0.955;0.957] [0.531;0.566] [0.946;0.948] [0.94;0.945] [0.946;0.948] [0.925;0.928] [0.926;0.93] [0.922;0.927] [0.893;0.897]
RF 0.96 0.161 0.951 0.978 0.951 0.932 0.935 0.929 0.903
[0.959;0.96] [0.155;0.168] [0.949;0.952] [0.977;0.979] [0.949;0.952] [0.93;0.934] [0.933;0.937] [0.926;0.931] [0.901;0.906]
ET 0.964 0.166 0.953 0.98 0.953 0.938 0.952 0.925 0.913
[0.963;0.965] [0.159;0.174] [0.952;0.954] [0.978;0.98] [0.952;0.954] [0.937;0.94] [0.95;0.954] [0.923;0.927] [0.911;0.915]
AdaBoost 0.888 0.574 0.856 0.894 0.856 0.806 0.838 0.776 0.729
[0.887;0.89] [0.574;0.575] [0.854;0.858] [0.892;0.897] [0.854;0.858] [0.803;0.809] [0.834;0.841] [0.772;0.78] [0.725;0.733]
GBM 0.895 0.244 0.867 0.916 0.867 0.82 0.845 0.795 0.747
[0.894;0.897] [0.242;0.246] [0.865;0.869] [0.914;0.918] [0.865;0.869] [0.817;0.822] [0.842;0.848] [0.792;0.799] [0.743;0.75]
LGBM 0.931 0.174 0.909 0.957 0.909 0.88 0.908 0.854 0.833
[0.93;0.932] [0.172;0.175] [0.907;0.91] [0.956;0.958] [0.907;0.91] [0.878;0.883] [0.906;0.911] [0.851;0.857] [0.83;0.836]
HistGB 0.929 0.175 0.907 0.956 0.907 0.877 0.902 0.853 0.827
[0.927;0.93] [0.173;0.176] [0.905;0.909] [0.955;0.957] [0.905;0.909] [0.875;0.879] [0.9;0.905] [0.85;0.856] [0.825;0.831]
XGBoost 0.928 0.172 0.912 0.956 0.912 0.878 0.885 0.872 0.827
[0.927;0.93] [0.17;0.174] [0.91;0.914] [0.955;0.957] [0.91;0.914] [0.877;0.881] [0.882;0.889] [0.869;0.875] [0.825;0.831]
XGBoost 0.883 - - - - 0.867 0.867 0.867 -
RF 0.881 - - - - 0.866 0.862 0.871 -
GBM 0.858 - - - - 0.844 0.836 0.856 -
SVC 0.848 - - - - 0.828 0.819 0.852 -
AdaBoost 0.840 - - - - 0.824 0.816 0.836 -
DT 0.834 - - - - 0.814 0.810 0.819 -
KNN 0.833 - - - - 0.814 0.809 0.820 -
LR 0.809 - - - - 0.796 0.789 0.822 -
Table 4. Performance comparison of the $\text{Chrome}_{\textbf{22}}$-trained classifiers. Best results are bold, second best are underlined, and baseline models are in italic. The performance results of the t.ex-Graph classifier are reported in the bottom half (Raschke et al., 2023).
Figure 5. The radar charts illustrate that while all models perform well on the training set, the XGBoost model exhibits more consistent performance across the various test sets, suggesting better generalizability. The other models show reduced performance on test sets, indicating potential overfitting to the training data.
In this section, we evaluate the effectiveness of our proposed methodology to detect trackers. We start by presenting our classifier results based on $\text{Chrome}_{22}$ data and follow-up by testing them with data from $\text{Firefox}_{22}$ and $\text{Brave}_{22}$. Furthermore, we assess the longitudinal performance with the $\text{Chrome}_{23}$ dataset that was collected almost a year later.
We test our approach on an Apple computer, namely MacBook Pro (14-inch, 2021) with a 10-core Central Processing Unit (CPU), a 16-core Graphical Processing Unit (GPU), 1 Terabyte (TB) of SSD storage, and 32 Gigabyte (GB) of Random Access Memory (RAM).
6.1. Intra-browser Performance
Table 4 shows the performance of our $\text{Chrome}_{\textbf{22}}$-trained classifiers and the baselines. In terms of Accuracy, Ensemble Trees (ET) lead with a score of 0.964, closely followed by Random Forest (RF) and Decision Trees (DT) with scores of 0.960 and 0.956, respectively.
The Log-Loss metric, which penalizes classifiers for incorrect predictions, shows that ET, RF, and DT also perform well, with ET having the lowest log loss, indicating a high level of confidence in its predictions. On the other hand, Gaussian Naive Bayes (GNB) exhibits a significantly higher log loss, which suggests that it is less confident in its predictions compared to the tree-based methods. The ROC-AUC metric further strengthens the position of ET, RF, and DT, as they all score above 0.94, indicating that they are good at differentiating between the positive and negative classes. The high MCC scores corroborate the findings from the other metrics, placing these three classifiers as the most robust performers across the board. In contrast, baseline models like LR are comparatively lower performance across these metrics.
The gradient boosting models, particularly XGBoost, have fairly high scores as reflected in an AUC-ROC score of 0.915 and an F1-Score of 0.875. They are able to handle the class imbalance due to the fairly high MCC and AUPRC scores. However, their precision and recall scores indicate room for improvement in identifying instances of the minority class.
Comparing the best performing classifiers ET and RF to the baseline models, shows that they perform significantly higher, except for the DT and its Log-Loss score. The t.ex-Graph classifiers display lower performance across all metrics.
6.2. Cross-browser and Longitudinal Performance
Figure 5 presents the cross-browser performance and longitudinal performance of our $\text{Chrome}_{\textbf{22}}$-trained classifiers. We focus on three selected classifiers, one baseline model, one ensemble tree-based model, and one ensemble tree-based model with gradient boosting.
For Logistic Regression and Random Forest, $\text{Chrome}_{\textbf{22}}$ shows high performance across most metrics, which is expected as the models would be most adapted to the data they were trained on. The performance on $\text{Chrome}_{\textbf{23}}$, $\text{Firefox}_{\textbf{22}}$, and $\text{Brave}_{\textbf{22}}$ varies, with a general trend of decreasing performance, indicating potential overfitting to the training data or a lack of generalizability to other datasets.
The XGBoost model shows a similar trend, with high performance on the training set and slightly varied results on the test sets. However, the variations in the XGBoost model’s performance are less pronounced than in the Logistic Regression and Random Forest models, suggesting that XGBoost may be more robust or generalizable across different browser datasets.
Overall, the figure indicates that while all models perform well on the dataset they were trained on, their ability to generalize to other data is varied. XGBoost appears to have the most consistent performance across different browser datasets, indicating a potential advantage in terms of generalization. The figure underscores the importance of cross-browser testing to ensure that models are not only tuned to a specific dataset but also maintain their performance across diverse scenarios.
The Brave results show a high accuracy metric for the overall performance, but the F1-Score and MCC are low and represent a more realistic performance of the classifiers, which is expected due to the more considerable differences between the Chrome and Brave data.
The longitudinal performance analysis, as depicted in Figure 5, indicates a slight decline in the efficacy of the classification models when transitioning from the $\text{Chrome}_{\textbf{22}}$ to the $\text{Chrome}_{\textbf{23}}$ dataset. This is evidenced by the contraction of the metric lines from $\text{Chrome}_{\textbf{22}}$ to $\text{Chrome}_{\textbf{23}}$ across the three models, suggesting a decrease in metrics such as AUC-ROC, Accuracy, and others. Notably, the XGBoost model demonstrates the smallest degree of performance reduction, underscoring its robustness and adaptability to changes over time. In contrast, both the Logistic Regression and Random Forest models exhibit a more pronounced drop, highlighting potential challenges in model generalization across browser version updates. This suggests that while all models are affected by evolving data characteristics, ensemble methods, particularly gradient boosting as implemented in XGBoost, may offer a more stable predictive performance in longitudinal settings. Applying isotonic calibration had only minor improvements compared to the uncalibrated models as shown in Figure 6.
7. Discussion
Tracking affects users across all three browsers. However, this varies to a certain degree, e.g. Brave had the least amount of tracking-related HTTP responses compared to Chrome and Firefox.
Nevertheless, an imbalance can be observed in all data sets, which might be a problem for specific ML models. Even though many different methods were developed to balance data sets, due to the moderate imbalance in Chrome and Firefox and the good performance, a resampling was not conducted to preserve the observed distribution. Still, the large number of observations provides sufficient HTTP response messages for each class label. Furthermore, several evaluation metrics were calculated and provided that are aimed at imbalanced data sets, e.g. BACC.
The classifiers performed similarly in Chrome and Firefox but were poor with Brave data. First and foremost, the application of binary-encoded features resulted in surprisingly good results, i.e. the mere presence of a header value is enough and extensive feature engineering is not required from a pure tracking detection standpoint. On the other hand, explaining the models is limited, except for model parameters, training time, or the calculated metrics. In addition, this method simplifies the models to a high degree, although HTTP header values can reveal more information about trackers and their differences from non-trackers. Nevertheless, the binary approach helped preserve the core information, namely whether a header is related to tracking.
The NB classifier did not perform as well as others, but that might be due to its limited applicability to the data sets. As mentioned above, NB is based on Bayes Theorem and assumes that features within a class are not correlated to each other. It can be assumed that a set of headers and thus features are correlated to a certain degree. These were not removed which is a violation of Bayes Theorem in a statistical sense could be argued, and thus, its applicability. On the contrary, a paper by Berrar (Berrar, 2018) argues that NB can perform well, even under these circumstances. Other implementations, e.g., the Categorical Naïve Bayes might have been a good fit for the training sets with categorical features (nb, 2023).
Filter lists are reliable and effective when it comes to blocking trackers (Mazel et al., 2019), and are one of the oldest methods, e.g. EasyList was created in 2005 (eas, 2023a). They present a popular ground truth to label data sets for training ML models. However, they have to be constantly updated, which commonly happens manually. A longitudinal study of nine years by Alrizah et al. (Alrizah et al., 2019) showed that human errors, e.g. acting in a timely manner or incorrect reports of possible new trackers and ads, are a root cause for lower performance. Chen et al. (Chen et al., 2021) highlighted evasion methods to circumvent blockers and filter lists: (i) changing URLs, (ii) embedding code directly into the website with the script tag, and (iii) combining functional and tracking code in one file. In summary, using filter lists for our ground truth is not without problems and directly affect the performance of our classifiers.
7.1. Limitations
Firstly, the analysis of response headers to fill the identified research gap has its strengths as well as its weaknesses. Headers represent only a part of an HTTP message, and research has shown that HTTP bodies are also relevant when it comes to understanding and identifying trackers (Li et al., 2015; Guarino. et al., 2020). From a broader point of view, only one part of application layer traffic is analyzed, disregarding other parts of web traffic, e.g. transport layer data.
Secondly, the data set could have been processed differently across the pipeline. As above-mentioned, the data sets are imbalanced, with a 72:28 non-tracker-to-tracker ratio for Chrome and Firefox data. It is not a severe form of imbalance, but it might be problematic for some classification algorithms. They could be biased against the majority class, and it might be possible to consistently predict the non-tracker label and still observe good metrics. Several resampling techniques exist, simple methods, e.g. over- and undersampling (Hernandez et al., 2013), or more complex methods, e.g. SMOTE (Chawla et al., 2002), or one of its variations (Han et al., 2005; Mukherjee and Khushi, 2021). Additional measures include using ML models, which handle imbalanced data sets well (Juez-Gil et al., 2021), or anomaly detection (Park et al., 2018). In summary, this aspect of the data pipeline has not been implemented with any of the aforementioned methods, and it might be interesting to explore how well these techniques may work for the present concept.
Thirdly, the labeling of the initial data sets, i.e. the quality of the ground truth. T.EX allows the implementation of different filter lists and offers three to date, namely EasyList, EasyPrivacy, and Disconnect.me. All base their functionality on extensive filter lists, i.e. rules and RegEx are used to block tracking-related HTTP requests. Updates are carried out regularly every four days for the first two (eas, 2023a, b), and on an irregular schedule for the latter (dis, 2023). Research has shown that these filter lists are not without errors regarding their rules (Snyder et al., 2020).
8. Conclusion
In this work, we trained Chrome-based classifiers on a subset of binary-encoded HTTP/S response headers to evaluate their effectiveness and usefulness in web tracker detection while retaining the natural imbalance of trackers. In addition, we measured the cross-browser effectiveness, showing that multiple Chrome-based classifiers achieve high accuracies for Firefox as well. However, large performance losses were observed for Brave. The longitudinal performance of the selected ML models deteriorated only by small margins, which highlights their potential due to the changing capabilities of trackers (Fouad et al., 2018; Bujlow et al., 2017; Englehardt and Narayanan, 2016). Response headers can be used to differentiate well between trackers and non-trackers and they provide additional information to the nature of tracking-related entities.
For our future research we plan to further investigate response headers, focusing on the values for feature engineering, or the use of Deep Learning models. Lastly, we will approach the deployment aspect two-fold, one from a user perspective in form of browser extension to test the models in a real-life scenario, and secondly as a fully automated system in the cloud that integrates a crawler to periodically update the models.